A printing machine valued at $3000 depreciates at a constant rate of 18% per year. In how many years will the machine be worth $750?


1
Expert's answer
2022-05-05T15:25:14-0400

"y=A(d)^x"

y=750

A=3000

d=1-0.18=0.82

x-?

"750=3000(0.82)^x"

"(0.82)^x=0.25"

"x=7 years"